OpenStudy (anonymous):

Find the domain: H(x)= sin^2radical x

OpenStudy (anonymous):

radical x?

OpenStudy (anonymous):

\[\sin ^{2}\sqrt{x}\]

OpenStudy (anonymous):

what do u understand by domain?

OpenStudy (anonymous):

domain is the set of all possible values of the function

OpenStudy (anonymous):

yea, but we look for what we CANT plug in can u think of anything?

OpenStudy (anonymous):

a negative number...?

OpenStudy (anonymous):

yes, so what will the domain be

OpenStudy (anonymous):

\[\ge0\] ?

OpenStudy (anonymous):

yes, u write it as \[[0,\infty)\] square bracket indicated 0 is included

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

can you help me to find the domain for \[\frac{ 1 }{ 1-\sin x }\]

OpenStudy (anonymous):

the denominator cant be zero